Explain if the limit of the function f(t) exist at t=1 where,

F(t)={ +5 if t<1} 

  {3-3t if t>1}


1
Expert's answer
2021-02-24T07:14:43-0500

If "\lim_{t\to1}f(t)" exists then the limits "\lim_{t\to1+0}f(t)" and "\lim_{t\to1-0}f(t)" must exist and be equal.

"\lim_{t\to1-0}f(t)=\lim_{t\to1-0}5=5"

"\lim_{t\to1+0}f(t)=\lim_{t\to1+0}(3-3t)=0"

Since these limits are distinct, the limit "\lim_{t\to1}f(t)" does not exist.